OpenStudy (anonymous):

find the value from 3 to 4 ((4x^2+2)/x^2) dx

OpenStudy (anonymous):

Integral?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

It had a dx after it, you should have inferred lol

OpenStudy (anonymous):

could have been a differential

OpenStudy (anonymous):

rewritten this way 4x^2/x^2 +2/x^2

OpenStudy (anonymous):

You can simplify it to 4 + 2/(x^2) dx and go from there

OpenStudy (anonymous):

4+2/x^2

OpenStudy (anonymous):

so do i plug in x as three and then solve. after that i add when i plug in x=4?

OpenStudy (anonymous):

\[\int\limits_3^4 \left(\frac{2}{x^2}+4\right) \, dx\]

OpenStudy (anonymous):

25/6

OpenStudy (anonymous):

im still confused. what im doing is just solving when x=3 and when x=4, then adding them, right?

OpenStudy (anonymous):

Did you find the integral first?

OpenStudy (anonymous):

You have to integrate 4 + 2/(x^2) before you plug in any values.

OpenStudy (anonymous):

thanks guys im all set!